OpenStudy (anonymous):

Explain, in complete sentences, which method you would use to solve the following system of equations and why you chose that method. Show your work and provide the solution to the system. 5x + 3y = –7 4x – 2y = –10

OpenStudy (ghazi):

(5x+3y=-7)*4 (4x-2y=-10)*5 now 20x+12y=-28 20x-10y=-50 now subtract both the equations 22y=22 y=1, put y=1 in first equation 5x=-7-3*1=-10 x=-2 and y=1...this is called as method of elimination